Name: Name Munir l) If air in a sealed flask st 20 Chas a pressure of 105 kPa. what is the value of k using the law of Gay-Lussac? 2) Use the value of k from question 1 to calculate the pressure of the air in the flask at
Answer #1

(1) T = 20 + 273 = 293 K

   P = 105 kPa

Now, 1 kPa = 0.00987 atm

So, 105 kPa = 105x0.00987 = 1.0364 atm

Again, as per Gay-Lussac law -

P / T = constant = k

=> k = 1.0364 / 293 = 0.00354atm/k = 3.54 x 10^-3 atm/k

(2) T = 273 + 40 = 313 K

Now -

P / T = k

=> P = k T

= 3.54 x 10^-3 x 313

= 1.108 atm